OpenStudy (anonymous):

How do you express the integral of e^(-x^(2)) as an infinite series??

OpenStudy (anonymous):

\[e^{-x^2}\] ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if so , \[e^x=\sum _{k=0}^{\infty } \frac{x^k}{k!}\] replace x with -x^2 \[\sum _{k=0}^{\infty } \frac{\left(-x^2\right)^k}{k!}\]

OpenStudy (anonymous):

That's all?? Oh wow xD

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

what about the integral part?

OpenStudy (anonymous):

I don't know about that

OpenStudy (zarkon):

integrate the series above term by term

OpenStudy (zarkon):

\[\int e^{-x^2}dx=\int\sum_{k=0}^{\infty}\frac{(-x^2)^k}{k!}dx=\sum_{k=0}^{\infty}\int\frac{(-x^2)^k}{k!}dx\]

OpenStudy (anonymous):

inf inf is sqrt pi...

OpenStudy (zarkon):

we don't have limits of integration

OpenStudy (zarkon):

\[\sum_{k=0}^{\infty}\int\frac{(-x^2)^k}{k!}dx=\sum_{k=0}^{\infty}\int\frac{(-1)^kx^{2k}}{k!}dx\] \[\sum_{k=0}^{\infty}\int\frac{(-x^2)^k}{k!}dx=\sum_{k=0}^{\infty}\frac{(-1)^k}{k!}\int x^{2k}dx\]

OpenStudy (anonymous):

I remeber doing this in a probability course..

OpenStudy (zarkon):

really....i doubt that...maybe this \[\int\limits_{-\infty}^{\infty}e^{-x^2}dx\] but that is a different problem than what we have here

OpenStudy (anonymous):

Yes, that one..root pi, right...

OpenStudy (zarkon):

sure...but the answer to the original problem is a function not a number

OpenStudy (anonymous):

You don't need limits of integrations of this. the final answer should be an integral of a series. leaving you with a final series but I figured it out anyways. It equals to the \[\sum_{n=0}^{\infty} ((-1)^{n}x ^{2n+1})/n!(2n+1)\]